Equipment and Calibration History
Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 [49] 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70Ser No. | Date | Loc From | Title | Loc To | Notes |
---|---|---|---|---|---|
NR-LITE 990352 | 2004-05-01 | Tonzi-Floor | Removed | Lab none | Remove net radiometer from met station near tower base - unit to the east of the tower |
NR-LITE 990352 | 2002-12-06 | Lab none | Installed | Tonzi-Floor | Installed on CR23x3 with the soil moisture probes at Tonzi. On the floor to the east of the tower. |
NR-LITE 990354 | 2002-06-06 | Lab none | Installed | Tonzi-Floor | Installed on CR23x3 with the soil moisture probes at Tonzi. |
CS451 70011229 | 2016-06-24 | Lab none | Installed | Mayberry Water Level | Installed on the Mayberry tower to replace the old sensor that has been acting flakey. |
CS451 20010046 | 2016-06-24 | Mayberry Water Level | Removed | Lab none | Removed from the Mayberry tower for testing. Replaced with new probe from DFW system |
CS451 70011236 | 2016-06-24 | Lab none | Installed | Twitchell Rice Water Level | Installed on the Twitchell Rice tower to replace the existing sensor that had a diurnal wave in its data. |
CS450 20010770 | 2016-06-24 | Twitchell Rice Water Level | Removed | Lab none | Removed from the Twitchell Rice tower due to diurnal wave in the data. Replace with new sensor from DFW equipment |
PTB110 H0160069 | 2016-06-24 | East End Barometer | Removed | Lab none | Removed from East End because it was reading low and erratically. |
CS511 DO-3 | 2015-09-02 | Mayberry DO Sensor C | Removed | Lab none | |
LI-820 CGA-556 | 2016-06-16 | Tonzi-Tower | Calibration | Tonzi-Tower | Last calibration was 2016-05-26 Cell T = 52.2, Pressure = 92.0 Zero Air offset = 5.0ppm Span CO2 405ppm offset = 0.0ppm Zero 0.9461 -> 0.9472 Span 0.8499 -> 0.8496 |
GP-001 GP20160003 | 2016-06-16 | Lab none | Installed | Tonzi-Tower | Installed at the 2nd lowest inlet to test as replacement for closed path system |
GP-001 GP20150084 | 2016-06-16 | Lab none | Installed | Tonzi-Tower | Installed at the lowest inlet to test as replacement for closed path system |
CR6 2115 | 2016-06-16 | Lab none | Installed | Tonzi-Tower | Installed to collect data from Eosense CO2 probes being tested for CO2 profile. |
BB-Fab BB-TF | 2016-06-16 | Tonzi-Floor | Upgrade | Tonzi-Floor | Set aside on battery that exploded. Other three tested okay. Added four more that were last at East End. |
BB-Fab BB-EE | 2016-03-24 | East End Battery Bank | Upgrade | East End Battery Bank | I swapped the batteries on the solar system. Four old batteries were removed and six new group 27 deep cycle flooded batteries were installed. Five of the new batteries have a date of 2/16 and one is 1/16. |
BB-Fab BB-EE | 2016-03-08 | East End Battery Bank | Maintenance | East End Battery Bank | I disconnected all the batteries and tested them individually. One of the older ones was warm to the touch and completely failed the load test. One of the batteries I added on 2015-11-20 also failed the load test. I removed these two batteries from the bank. |
LI-7500A 75H-2514 | 2016-06-03 | Lab none | Calibration | Lab none | Last calibration: 2016-02-09 AGC=103.3 SDM=6 DAC1=- DAC2=- Ambient P=100.04kPa, Tair=25.86C, Tdew=13.67C Zero Air at 500psi, God CO2 421.22 at 500 Tdewspan=15.08 CO2 zero 0.905 -> 0.9047 Offset: -1.32ppm CO2 span 1.0028 -> 1.0055 Offset: -1.52ppm H2O zero 0.8364 -> 0.8368 Offset: 15.3ppm H2O span 1.0226 -> 1.0150 Offset: 0.16ppm Notes: The spars were very dirty with clinging dust. |
LI-7500A 75H-2176 | 2016-06-03 | Lab none | Calibration | Lab none | Last calibration: 2016-02-22 AGC=90.7 SDM=6 DAC1=- DAC2=- Ambient P=100.05kPa, Tair=24.56C, Tdew=12.80C Zero Air at 500psi, God CO2 421.22 at 500 Tdewspan=15.08 CO2 zero 0.8692 -> 0.8688 Offset: -1.68ppm CO2 span 1.0205 -> 1.0175 Offset: 1.63ppm H2O zero 0.8564 -> 0.8571 Offset: 4.3ppm H2O span 1.0136 -> 1.0281 Offset: -0.3ppm Notes: The spars were very dirty with clinging dust. |
LI-7500A 75H-2514 | 2016-06-06 | Lab none | Installed | Twitchell Alfalfa Gas Analyzer CO2/H2O | Installed at Twitchell alfalfa to replace sn 75H-2180 for calibration. |
LI-7500A 75H-2176 | 2016-06-06 | Lab none | Installed | West Pond Gas Analyzer CO2/H2O | installed at west pond to replace sn 1594 for calibration. |
LI-7500A 1594 | 2016-06-06 | West Pond Gas Analyzer CO2/H2O | Removed | Lab none | removed from west pond for calibration, replaced by sn 75H-2176 |
LI-7500A 75H-2180 | 2016-06-06 | Twitchell Alfalfa Gas Analyzer CO2/H2O | Removed | Lab none | Removed from tower in Twitchell alfalfa for calibration, replaced by sn 75H-2514. |
TR-525I-R1 24875-999 | 2010-09-10 | Lab none | Installed | Vaira Precipitation | Reinstalled on the Vaira tower after testing in the lab. The testing revealed that the problem was a bug in the data logger program that was recording instantaneous values instead of 30min averages. |
KU265-6MCA 158APW4M0000190 | 2016-05-24 | Lab 0 | Installed | Sherman Wetland Solar Panel A | Installed on the Sherman Wetland tower at the initial set up of the site. |
KU265-6MCA 158APW4M0000193 | 2016-05-24 | Lab 0 | Installed | Sherman Wetland Solar Panel B | Installed on the Sherman Wetland tower at the initial set up of the site. |
SD130BN 0030F4D21491 | 2016-06-06 | Lab none | Installed | Sherman Wetland Camera | Initial installation at the Sherman Wetland Tower. |
GP-001 GP20150101 | 2016-05-24 | Lab none | Installed | Sherman Wetland CO2 Probe | Installed on the Sherman Wetland tower at the initial set up of the site. In the air unil the site is flooded. |
PQS 1 140454 | 2016-05-24 | Lab none | Installed | Sherman Wetland PAR Sensor Incoming | Installed on the Sherman Wetland tower at the initial set up of the site. It is in the incoming posistion. |
PQS 1 140450 | 2016-05-24 | Lab none | Installed | Sherman Wetland PAR Sensor Reflected | Installed on the Sherman Wetland tower at the initial set up of the site. It is in the reflected posistion. |
PTB110 L2510774 | 2016-05-24 | Lab none | Installed | Sherman Wetland Barometer | Installed on the Sherman Wetland tower at the initial set up of the site. Installed in the data logger box. |
HMP60 L2310736 | 2016-05-24 | Lab none | Installed | Sherman Wetland RH & Temp | Installed on the Sherman Wetland tower at the initial set up of the site. |
HFP01 11325 | 2016-05-24 | Lab none | Installed | Sherman Wetland Heat Flux Plate C | Installed on the Sherman Wetland tower at the initial set up of the site. |
HFP01 11329 | 2016-05-24 | Lab none | Installed | Sherman Wetland Heat Flux Plate B | Installed on the Sherman Wetland tower at the initial set up of the site. |
HFP01 11328 | 2016-05-24 | Lab none | Installed | Sherman Wetland Heat Flux Plate A | Installed on the Sherman Wetland tower at the initial set up of the site. |
SRS-Ni NDVI 882103059 | 2016-05-24 | Lab none | Installed | Sherman Wetland Narrow Band NDVI Incoming | Installed on the Sherman Wetland tower at the initial set up of the site. It is in the incoming posistion and using SDM address C. |
SRS-Pi PRI 912903006 | 2016-05-24 | Lab none | Installed | Sherman Wetland Narrow Band PRI Incoming | Installed on the Sherman Wetland tower at the initial set up of the site. It is in the incoming posistion and using SDM address A. |
SRS-Nr NDVI 943203022 | 2016-05-24 | Lab none | Installed | Sherman Wetland Narrow Band NDVI Reflected | Installed on the Sherman Wetland tower at the initial set up of the site. It is in the reflected posistion and using SDM address D. |
SRS-Pr PRI 930203026 | 2016-05-24 | Lab none | Installed | Sherman Wetland Narrow Band PRI Reflected | Installed on the Sherman Wetland tower at the initial set up of the site. It is in the reflected posistion and using SDM address B. |
CS451 70011234 | 2016-05-24 | Lab none | Installed | Sherman Wetland Water Level | Installed on the Sherman Wetland tower at the initial set up of the site. |
AM25T 7909 | 2016-05-24 | Lab none | Installed | Sherman Wetland Multiplexer Thermocouple | Installed on the Sherman Wetland tower at the initial set up of the site. |
NR01 2466 | 2016-05-24 | Lab none | Installed | Sherman Wetland Net Radiometer 4-way | Installed on the Sherman Wetland tower at the initial set up of the site. |
TE525MM 65552-615 | 2016-05-24 | Lab none | Installed | Sherman Wetland Precipitation | Installed on the Sherman Wetland tower at the initial set up of the site. |
AM16/32B 21689 | 2016-05-24 | Lab none | Installed | Sherman Wetland Multiplexer Met | Installed on the Sherman Wetland tower at the initial set up of the site. |
CR1000 38211 | 2016-05-24 | Lab none | Installed | Sherman Wetland Data Logger Met | Installed on the Sherman Wetland tower at the initial set up of the site. |
WM 1590 153803 | 2016-05-24 | Lab none | Installed | Sherman Wetland Anemometer | Installed on the Sherman Wetland tower at the initial set up of the site. |
LI-7550 AIU-1569 | 2016-05-24 | Lab none | Installed | Sherman Wetland Interface LI-7550 | Installed on the Sherman Wetland tower at the initial set up of the site. |
LI-7700 TG1-0423 | 2016-05-24 | Lab none | Installed | Sherman Wetland Gas Analyzer CH4 | Installed on the Sherman Wetland tower at the initial set up of the site. |
LI-7500A 75H-2668 | 2016-05-24 | Lab none | Installed | Sherman Wetland Gas Analyzer CO2/H2O | Installed on the Sherman Wetland tower at the initial set up of the site. |
NL115 13054 | 2016-06-02 | Lab none | Installed | Sherman Wetland Interface NL115 | Added to CR1000 sn 38211 at the Sherman Wetland tower in preparation of installing a Stardot Phenocam |
LI-820 CGA-556 | 2016-05-26 | Tonzi-Tower | Calibration | Tonzi-Tower | Last calibration was 2016-02-09 Cell T = 52.2, Pressure = 92.0 Zero Air offset = 16.4ppm Span CO2 405ppm offset = -19.4ppm Zero 0.9423 -> 0.9461 Span 0.8220 -> 0.8499 Changed Pressure output range from 1.33331998E3 to 115kPa |
LI-7500A 0065 | 2016-05-24 | Lab none | Installed | Twitchell Rice Gas Analyzer CO2/H2O | Installed at Twitchell Rice tower to replace sn 75H-2176 for calibration |
LI-7500A 75H-2176 | 2016-05-24 | Twitchell Rice Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Twitchell Rice tower for calibration and replaced with sn 0065 |
LI-7500A 75H-2150 | 2016-05-24 | Lab none | Installed | East End Gas Analyzer CO2/H2O | Installed LI7500 at East End tower, replaced sn 75H-2514 which was taken down for calibration. |
LI-7500A 75H-2514 | 2016-05-24 | East End Gas Analyzer CO2/H2O | Removed | Lab none | Removed LI7500 from East End tower for calibration. Replaced with LI 75H-2150. |
LI-7500A 75H-2668 | 2016-05-20 | Lab none | Upgrade | Lab none | Upgraded embedded software from 7.5.0 to 8.0.0 |
LI-7700 TG1-0423 | 2016-05-20 | Lab none | Calibration | Lab none | Calibration Location: Lab Last calibration: 2015-07-23 (Factory) RSSI=79.4%, Optical RH=18% Ambient P=100.08kPa, Tair=22.6C Zero Air at 500psi, 17-07-2013 CH4 Tank 4.995 at 500 CH4 Zero Offset: 0.03ppm CH4 Span Offset: 0.006ppm Notes: Uneventful calibration. Previous calibration was factory. |
LI-7500A 75H-2668 | 2016-05-20 | Lab none | Calibration | Lab none | Last calibration: 2015-07-22 (Factory) AGC=100.22 SDM= DAC1= DAC2= Ambient P=100.03kPa, Tair=21.23C, Tdew=9.06C Zero Air at 600psi, NOAA "Good Co2" 421.22 at 500 Tdewspan=15.09 CO2 zero 0.9380 -> 0.9385 Offset: 1.79ppm CO2 span 1.0021 -> 1.0136 Offset: 6.22ppm H2O zero 0.8966 -> 0.8969 Offset: 1.4ppm H2O span 1.0037 -> 1.0146 Offset: 0.23ppm Notes: Upgraded from 7.5.0 to 8.0.0 embedded software. Last calibration was Factory - all setting from Factory cal were saved. Factory cal done at 749.7 ppm and 11.98 degC. |
SD130BN 0030F4D2148D | 2016-05-17 | Lab none | Installed | Twitchell Alfalfa | Installed on the camera post to over lap with existing Canon camera |
ML2x 377/008 | 2016-05-17 | Lab none | Installed | Twitchell Alfalfa | Installed at -10cm to the southeast of the soil co2 box |
ML2x 377/010 | 2016-05-12 | Lab none | Installed | Twitchell Alfalfa | Installed at -20cm to the northeast of the soil co2 box |
HFP01 6491 | 2016-05-17 | Twitchell Alfalfa | Moved | Twitchell Alfalfa | moved from 3rd (west) channel on met CR1000 to the 3rd (NE) channel on the soil co2 CR1000 |
HFP01 6498 | 2016-05-17 | Twitchell Alfalfa | Moved | Twitchell Alfalfa | moved from 2nd (NW) channel on met CR1000 to the 2nd (SE) channel on the soil co2 CR1000 |
HFP01 6342 | 2016-05-12 | Twitchell Alfalfa | Moved | Twitchell Alfalfa | moved from 1st (W) channel on met CR1000 to the 1st (NE) channel on the soil co2 CR1000 |
LI-7500A 0035 | 2016-05-10 | Lab none | Installed | Tonzi-Floor | Installed on the Tonzi Floor to replace sn 0042 for calibration. |
LI-7500A 0042 | 2016-05-10 | Tonzi-Floor | Removed | Lab none | Removed from the Tonzi Floor for calibration. Replaced by sn 0035. |
LI-7500A 0041 | 2016-05-10 | Lab none | Installed | Tonzi-Tower | Installed on the Tonzi Tower to replace sn 0418 for calibration. |
LI-7500A 0418 | 2016-05-10 | Tonzi-Tower | Removed | Lab none | Removed from the Tonzi Tower for calibration. Replaced by sn 0041. |
LI-7500A 0073 | 2016-05-10 | Lab none | Installed | Vaira Gas Analyzer CO2/H2O | Installed on the Vaira tower to replace sn 0976 for calibration. |
LI-7500 0976 | 2016-05-10 | Vaira Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Vaira tower for calibration. Replaced by sn 0073. |
LI-7500A 0065 | 2016-05-13 | Lab none | Calibration | Lab none | Last calibration: 2015-12-07 AGC=74.96 SDM=- DAC1=- DAC2=- Ambient P=100.14kPa, Tair=20.14C, Tdew=8.2C Zero Air at 600psi, God CO2 421.22 at 500 Tdewspan=15.09 CO2 zero 0.89 -> 0.8898 Offset: -0.57ppm CO2 span 1.0164 -> 1.0149 Offset: 0.91ppm H2O zero 0.7097 -> 0.7086 Offset: -7.4ppm H2O span 1.0231 -> 1.0139 Offset: 0.19ppm Notes: Calibration from H2O performed at 15.09C dew point instead of 11.09 which has been used previously. Use 15.09 from now on for a larger calibrated H2O range. |
LI-7500A 75H-2150 | 2016-05-13 | Lab none | Calibration | Lab none | Last calibration: 2015-12-07 AGC=97.6 SDM=- DAC1=- DAC2=- Ambient P=100.15kPa, Tair=20.52C, Tdew=8.5C Zero Air at 600psi, God CO2 421.22 at 500 Tdewspan=15.09 CO2 zero 0.9019 -> 0.9020 Offset: 0.40ppm CO2 span 1.0114 -> 1.0080 Offset: 1.84ppm H2O zero 0.8671 -> 0.8668 Offset: -2.4ppm H2O span 1.0258 -> 1.0109 Offset: 0.31ppm Notes: Calibration from H2O performed at 15.09C dew point instead of 11.09 which has been used previously. Use 15.09 from now on for a larger calibrated H2O range. |
LI-7500A 0035 | 2016-05-05 | Lab none | Calibration | Lab none | Last calibration: 2015-10-28 AGC=49 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=100.0kPa, Tair=19.34C, Tdew=8.30C Zero Air at 600psi, God CO2 421.22 at 500 Tdewspan=11.09 CO2 zero 0.9384 -> 0.9388 Offset: 1.46ppm CO2 span 1.0237 -> 1.0216 Offset: 1.08ppm H2O zero 0.903 -> 0.9018 Offset: -5.9ppm H2O span 0.8731 -> 0.9928 Offset: -2.5ppm Notes: Large offset in H2O span. I think a mistake was made last time when the dew point was set to the H2O mmol/mol value. |
LI-7500A 0041 | 2016-05-05 | Lab none | Calibration | Lab none | Last calibration: 2015-10-28 AGC=50 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=100.03kPa, Tair=19.99C, Tdew=9.10C Zero Air at 600psi, God CO2 421.22 at 500 Tdewspan=11.09 CO2 zero 1.0004 -> 1.0002 Offset: -0.7ppm CO2 span 1.0107 -> 1.0128 Offset: -1.12ppm H2O zero 0.8247 -> 0.8257 Offset: 5.0ppm H2O span 0.8836 -> 1.0037 Offset: -2.5ppm Notes: Large offset in H2O span. I think a mistake was made last time when the dew point was set to the H2O mmol/mol value. |
LI-7500A 0073 | 2016-05-05 | Lab none | Calibration | Lab none | Last calibration: 2015-10-28 AGC=58 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=99.90kPa, Tair=19.2C, Tdew=8.98C Zero Air at 600psi, God CO2 421.22 at 500 Tdewspan=11.09 CO2 zero 0.9584 -> 0.9579 Offset: -1.8ppm CO2 span 1.0034 -> 1.0029 Offset: 0.18ppm H2O zero 0.7403 -> 0.7385 Offset: -9.8ppm H2O span 0.9201 -> 1.0159 Offset: -2.01ppm Notes: Large offset in H2O span. I think a mistake was made last time when the dew point was set to the H2O mmol/mol value. |
CS547A 5688 | 2016-04-28 | Lab none | Installed | Mayberry Conductivity Sensor | Installed new conductivity sensor. The old one was removed last time since it was reading weird numbers. |
GP-001 GP20140084 | 2016-04-28 | Lab none | Installed | Twitchell Alfalfa | Replaced probe sn GP20140023 at 2cm depth since that was reading too high. |
GP-001 GP20140023 | 2016-04-28 | Twitchell Alfalfa | Removed | Lab none | Probe was reading really high numbers. Brought back to lab for repairs. Replaced with sn 20140084 |
NOAA-CMDL CA07091 | 2013-01-23 | Factory | Disposed | Factory | This tank was traded or NOAA-ERL in exchange for a new tank sn CB09920 |
NOAA-CMDL CA07091 | 2013-01-23 | Factory | Calibration | Factory | Post use NOAA_ERL calibration 1898.51ppb CH4 |
NOAA-CMDL CA07091 | 2012-12-14 | Lab none | Shipped | Factory | This tank was shipped to NOAA_ERL for a post use calibration. It will be retained by NOAA in exchange for a new tank. |
NOAA-CMDL CA04396 | 1999-08-01 | Factory | Tested | Factory | Original hydrostatic test 08A99 (August 1999). Next due August 2004. |
NOAA-CMDL CA04119 | 1999-06-01 | Factory | Tested | Factory | Original hydrostatic test 06A99 (June 1999). Next due June 2004. |
NOAA-CMDL CA03923 | 1999-02-01 | Factory | Tested | Factory | Original hydrostatic test 02A99 (February 1999). Next due February 2004. |
NOAA-CMDL CA03911 | 1999-02-01 | Factory | Tested | Factory | Original hydrostatic test 02A99 (February 1999). Next due February 2004. |
NOAA-CMDL CA04951 | 2001-03-01 | Factory | Tested | Factory | Original hydrostatic test 03A01 (March 2001). Next due March 2006. |
NOAA-CMDL CB09920 | 2012-06-01 | Factory | Tested | Factory | Original hydrostatic test 06A12 (June 2012). Next due June 2017. |
LI-7550 AIU-0614 | 2016-04-11 | Lab none | Upgrade | Lab none | Updated the 7550 firmware from v7.3.0 to v8.0.0 skiping version 7.5.0 Relevant 7.5.0 changes: Main Window is now resizable, Added serial number entry for sonics, Help update, Fixed timeout when saving config files Relevant 8.0.0 changes: Separate bandwidth settings for analog inputs and gas channels, Detection of new RS instruments, Additional calibration coefficients, Added sonic software version entry, Added chopper temp logging |
LI-7500A 75H-2182 | 2016-04-11 | Lab none | Calibration | Lab none | Last calibration: 2015-11-10 AGC=87.6 SDM=6 DAC1=- DAC2=- Ambient P=100.48kPa, Tair=18.11C, Tdew=9.67C Zero Air at 600psi, God CO2 421.22 at 500 Tdewspan=11.09 CO2 zero 0.8833 -> 0.8832 Offset: -0.23ppm CO2 span 1.0197 -> 1.0198 Offset: -0.02ppm H2O zero 0.8464 -> 0.8466 Offset: 0.8ppm H2O span 1.0242 -> 1.0138 Offset: 0.21ppm Notes: Firmware on AIU-0614 updated from v7.3.0 to v8.0.0 just before calibration |
LI-7550 AIU-1285 | 2016-04-13 | Mayberry Interface LI-7550 | Upgrade | Mayberry Interface LI-7550 | Updated the 7550 firmware from v7.3.0 to v8.0.0 skiping version 7.5.0 Relevant 7.5.0 changes: Main Window is now resizable, Added serial number entry for sonics, Help update, Fixed timeout when saving config files Relevant 8.0.0 changes: Separate bandwidth settings for analog inputs and gas channels, Detection of new RS instruments, Additional calibration coefficients, Added sonic software version entry, Added chopper temp logging |
A547 3270 | 2016-04-13 | Mayberry Interface A547 | Removed | Lab none | Removed from Mayberry with conductivity probe sn 6097 due to bad readings. Needs to be tested in the lab. |
LI-7550 AIU-0674 | 2016-04-13 | Twitchell Alfalfa Interface LI-7550 | Upgrade | Twitchell Alfalfa Interface LI-7550 | Updated the 7550 firmware from v7.3.0 to v8.0.0 skiping version 7.5.0 Relevant 7.5.0 changes: Main Window is now resizable, Added serial number entry for sonics, Help update, Fixed timeout when saving config files Relevant 8.0.0 changes: Separate bandwidth settings for analog inputs and gas channels, Detection of new RS instruments, Additional calibration coefficients, Added sonic software version entry, Added chopper temp logging |
LI-7550 AIU-0666 | 2016-04-13 | West Pond Interface LI-7550 | Upgrade | West Pond Interface LI-7550 | Updated the 7550 firmware from v7.3.0 to v8.0.0 skiping version 7.5.0 Relevant 7.5.0 changes: Main Window is now resizable, Added serial number entry for sonics, Help update, Fixed timeout when saving config files Relevant 8.0.0 changes: Separate bandwidth settings for analog inputs and gas channels, Detection of new RS instruments, Additional calibration coefficients, Added sonic software version entry, Added chopper temp logging |
LI-7550 AIU-0397 | 2016-04-13 | East End Interface LI-7550 | Upgrade | East End Interface LI-7550 | Updated the 7550 firmware from v7.3.0 to v8.0.0 skiping version 7.5.0 Relevant 7.5.0 changes: Main Window is now resizable, Added serial number entry for sonics, Help update, Fixed timeout when saving config files Relevant 8.0.0 changes: Separate bandwidth settings for analog inputs and gas channels, Detection of new RS instruments, Additional calibration coefficients, Added sonic software version entry, Added chopper temp logging. |
CS511 DO-8 | 2016-04-13 | Twitchell Rice DO Sensor | Removed | Lab none | Removed with tower from field for planting. Needs to be serviced in the lab and reinstalled when the field is flooded. |
LI-7550 AIU-0650 | 2016-04-13 | Twitchell Rice Interface LI-7500 | Upgrade | Twitchell Rice Interface LI-7500 | Updated the 7550 firmware from v7.3.0 to v8.0.0 skiping version 7.5.0 Relevant 7.5.0 changes: Main Window is now resizable, Added serial number entry for sonics, Help update, Fixed timeout when saving config files Relevant 8.0.0 changes: Separate bandwidth settings for analog inputs and gas channels, Detection of new RS instruments, Additional calibration coefficients, Added sonic software version entry, Added chopper temp logging |
LI-7700 TG1-0155 | 2016-04-13 | East End Gas Analyzer CH4 | Calibration | East End Gas Analyzer CH4 | Calibration Location: East End Last calibration: 2015-09-30 RSSI=86.86%, Optical RH=12% Ambient P=101.2kPa, Tair=20.94C Zero Air at 950psi, AMP CH4 1.9966 at 850 CH4 Zero Offset: 0.002ppm CH4 Span Offset: -0.010ppm Notes: Gusty breeze |
LI-7700 TG1-0223 | 2016-04-13 | West Pond Gas Analyzer CH4 | Calibration | West Pond Gas Analyzer CH4 | Calibration Location: West Pond Last calibration: 2015-09-30 RSSI=84.2%, Optical RH=24% Ambient P=101.74kPa, Tair=22.97C Zero Air at 900psi, AMP CH4 1.9966 at 700 CH4 Zero Offset: 0.011ppm CH4 Span Offset: 0.001ppm Notes: |
LI-7700 TG1-0211 | 2016-04-13 | Twitchell Rice Gas Analyzer CH4 | Calibration | Twitchell Rice Gas Analyzer CH4 | Calibration Location: Twitchell Rice Last calibration: 2015-09-30 RSSI=66.1%, Optical RH=23% Ambient P=102.1kPa, Tair=18.6C Zero Air at 1000psi, AMP CH4 1.9966 at 950 CH4 Zero Offset: 0.018ppm CH4 Span Offset: -0.010ppm Notes: Gusty breeze, mirrors hard to get clean. |
LI-7700 TG1-0357 | 2016-04-13 | Mayberry Gas Analyzer CH4 | Calibration | Mayberry Gas Analyzer CH4 | Calibration Location: Mayberry
Last calibration: 2015-09-30 RSSI=78.3%, Optical RH=7.5% Ambient P=102.1kPa, Tair=15.8C Zero Air at 1000psi, AMP CH4 1200 at 1.9966 CH4 Zero Offset: -0.024ppm CH4 Span Offset: -0.009ppm Notes: Gusty breeze, dropped the hood in the water before calibration - dried it off. |
CS547A 6097 | 2016-04-13 | Mayberry Conductivity Sensor | Removed | Lab none | Removed from Mayberry tower together with interface module since it was giving bad readings. |